6(x-7)-2(3+4)+3(x)=34

Answer by checkley71(8403)

6(X-7)-2(3+4)+3(X)=34
6X-42-2*7+3X=34
6X+3X-42-14=34
9X=34+42+14
9X=90
X=90/9
X=10 ANSWER.
PROOF
6(10-7)-2*7+3*10=34
6*3-14+30=34
18-14+30=34
48-14=34
34=34